OpenStudy (anonymous):

Which tables show a direct variation? Choose all answers that are correct. x y 1 1.2 2 2.4 3 3.6 A. x y 5 15 10 30 15 60 B. x y 1 4 3 6 5 8 C. x y 3 6 6 12 9 18 D.

OpenStudy (anonymous):

what is the rule for direct variation ?

OpenStudy (anonymous):

and what is option D ?

OpenStudy (anonymous):

oh there going up

OpenStudy (anonymous):

ohh... hmm alright then... how do you see if a table has direct variation or not ?

OpenStudy (anonymous):

dude are u hacking

OpenStudy (anonymous):

nope... I am not...

OpenStudy (anonymous):

what the heck

OpenStudy (anonymous):

it is still there

OpenStudy (anonymous):

lets answer the question that you have posted... I dont know why yo are seeing so many medals there

OpenStudy (anonymous):

ok lets do it

OpenStudy (anonymous):

*you

OpenStudy (anonymous):

i dont hack

OpenStudy (anonymous):

the last response was for the typo :) now, how do you see if a table has direct variation or not ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

my last response :)

OpenStudy (anonymous):

anyhow... how do you see if a table has direct variation or not ?

OpenStudy (anonymous):

no i dont know this

OpenStudy (anonymous):

it is a new school

OpenStudy (anonymous):

any table or set of values that satisfy y=k*x where k is constant is said to have direct variation see which of the above options fit the criterion...

OpenStudy (anonymous):

Choose all answers that are correct. x y 1 1.2 2 2.4 3 3.6 A. x y 5 15 10 30 15 60 B. x y 1 4 3 6 5 8 C. x y 3 6 6 12 9 18 D.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

lets take the first one as example... 1.2 = (1.2)*1 2.4 = (1.2)*2 3.6=(1.2)*3 here y=k*x where the constant k=1.2... hence option A has direct variation... can you try the remaining ones and let me know ?

OpenStudy (anonymous):

sup jhonyy

OpenStudy (anonymous):

the all are devided the same

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

option B does not have direct variation... can you tell me why ??

OpenStudy (anonymous):

its what i got on the question

OpenStudy (anonymous):

lets look at option B: if it had direct variation y=k*x 15=(3)*5 => k=3 30=(3)*10 => k=3 60=(4)*15 => k=4 as the value of k is different for one of the pairs... the table does not have direct variation... Now can you try the same with remaining options ??

OpenStudy (anonymous):

@MLGNAME